OpenStudy (anonymous):

Which is the equation of a circle with the center at (4, 7) and the radius 4 units long? (x + 4)2 + (y + 7)2 = 16 (x − 4)2 + (y − 7)2 = 16 (x + 4)2 + (y + 7)2 = 4 (x − 4)2 + (y − 7)2 = 4

OpenStudy (anonymous):

(x-4)2+(y-7)2=16

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

the general equation of a circle is (x-a)^2+(y-b)^2=r^2 when (a,b) is the center.r is radius